# Dynamic Rank Adjustment for Accurate and Efficient Neural Network Training

### v2

@@ Abstract
Low-rank training methods reduce the number of trainable parameters by re-parameterizing the weights with matrix decompositions (e.g., singular value decomposition). However, enforcing a fixed low-rank structure caps the rank of the weight matrices and can hinder the model’s ability to learn complex patterns. Furthermore, the effective rank of the model’s weights tends to decline during training, and this drop is accelerated when the model is reparameterized into a low-rank structure. In this study, we argue that strategically interleaving full-rank training epochs within low-rank training epochs can effectively restore the rank of the model’s weights. Based on our findings, we propose a general dynamic-rank training framework that is readily applicable to a wide range of neural-network tasks. We first describe how to adjust the rank of weight matrix to alleviate the inevitable rank collapse that arises during training, and then present extensive empirical results that validate our claims and demonstrate the efficacy of the proposed framework. Our empirical study shows that the proposed method achieves almost the same computational cost as SVD-based low-rank training while achieving a comparable accuracy to full-rank training across various benchmarks.
@@ 1 Introduction
Low-rank reparameterization methods have been actively studied for efficient training of large neural networks. Low-rank training strategies typically reduce the number of trainable parameters by applying matrix decompositions to a model’s weight matrices. While this approach lowers training cost, it permanently caps the maximum attainable rank of those matrices, thereby limiting the model’s ability to learn complex patterns. Moreover, recent studies report that the effective rank of the weights tends to decline during training [32, 11]. Therefore, to train large neural networks efficiently while preserving their learning capacity, it is crucial to address the decline in weight-matrix rank that occurs during low-rank training.
@@
Singular value decomposition (SVD)–based method [14] is among the most widely used re-parameterization techniques. Other tensor factorizations, such as Tucker [18] and CP [20] decompositions, have likewise been adopted to enable low-rank training. In federated-learning settings, FLANC [23] and FedPara [12] re-parameterize models into low-rank forms. Although these approaches reduce the number of trainable parameters, they often ignore the progressive decline in the weights’ effective rank. Some regularization methods, such as soft orthogonal regularizer [32] and its variants [1, 17], focus on the rank decline issue. However, they incur higher computational cost and memory overhead, making them less practical for training large neural networks.
@@
This study explores how to mitigate the decline in the effective rank of model weights in low-rank training. Our key finding is that interleaving a few full-rank epochs within low-rank training effectively preserves the model’s effective rank. Specifically, we analyze how run-time adjustments to the rank of model weights affect training and present a practical strategy for raising or lowering rank during low-rank training to maximize the model’s effective rank. Our theoretical analysis and empirical study demonstrate that dynamic rank adjustment matches the accuracy of full-rank training while retaining the system efficiency of low-rank methods.
@@
Another key finding is that the effectiveness of rank adjustment is closely coupled with the learning rate. Our study shows that full-rank epochs should be scheduled according to the noise scale induced by the learning rate. In general, the learning rate tends to be initialized to a large value and then progressively decayed during training. In this study, we investigate how the learning rate influences the gap between low-rank and full-rank updates and, based on this analysis, develop a general rank scheduling framework that maximizes the benefits of interleaving full-rank epochs within low-rank training epochs. Table 1 shows a feature-wise comparison among various low-rank training methods.
@@
To evaluate the performance of our proposed dynamic-rank training framework, we conduct an extensive empirical study on widely used computer vision and natural language processing benchmarks. We further benchmark our approach against state-of-the-art (SOTA) low-rank training techniques and regularization methods designed to restore effective rank. Throughout all the experiments, we find that our dynamic-rank training method achieves comparable accuracies to the full-rank training while significantly reducing the computational cost like conventional low-rank training methods. In addition, our experiments demonstrate that the proposed method integrates seamlessly with low-rank training when paired with soft-orthogonality (SO) regularization, confirming that the two techniques are complementary. we therefore conclude that the dynamic rank adjustment is a practical and easy-to-use training strategy that effectively boosts model accuracy in low-rank training.
@@ 2 Background
Method PE $R$ PC NA
Low-rank (SVD, etc.)
GKPD
PELA
TKD-CPD
Soft Orthogonality (SO)
SRIP
Dynamic-rank
Table 1: A feature-wise comparison between dynamic-rank training and other low-rank training methods. PE: parameter-efficient, $R$: rank recovery, PC: pre-training compatible, and NA: noise-scale aware.
Low-rank Re-parameterization – Low-rank reparameterization is a model approximation technique that is popularly used in large neural network training. Given a model weight matrix $\mathbf{W}$, a matrix decomposition method is applied to $\mathbf{W}$ before training begins. For example, if SVD is used, $\mathbf{W}\in\mathbb{R}^{m\times n}$ is decomposed to two smaller matrices $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ such that $\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$, where $k<m$ and $k<n$. As $k$ decreases, the total number of model parameters is reduced, thereby lowering the computational cost of training. However, the reconstructed weight matrix can have up to $k$ ranks, resulting in the limited learning capability. Low-rank Adaptation (LoRA) [10] is one of the most popular applications of re-parameterization technique, particularly in the context of fine-tuning. The importance of singular values under LoRA’s context is deeply analyzed in  [16].
Learning Rate and Noise Scale – In previous works, learning rate is known to play a key role in determining generalization performance of machine learning models. It has been theoretically shown that the noise scale $g$ in gradient approximation is determined by the learning rate and batch size, as follows [27]: $g\approx\eta N/B$, where $\eta$ is learning rate, $N$ is the dataset size, and $B$ is the mini-batch size. This relation links learning rate schedules to the notion of noise scale. In early training, a large noise scale helps the model explore and shape the decision boundary [22, 21], consistent with the common practice of decaying the learning rate. In this study, we investigate the relationship between the noise scale and the effective rank of the model’s weights.
Figure 1: Comparison of the layer-wise singular-value spectral ratio ($\lambda$) across different model ranks during ResNet20 training on CIFAR-10. The left plot shows layer-wise $\lambda$ curves for the full-rank model training while the right plot shows the $\lambda$ curves for the reparameterized model training. All convolution layers were reparameterized to have half rank. At the end of the training, many layers in the reparameterized model have large $\lambda$ values, meaning the model has converges to a low rank space.
@@ 3 Related Work
Low-rank Reparameterization Methods – Several studies have explored the use of low-rank re-parameterization techniques including singular value decomposition, Tucker decomposition, and CP decomposition methods [14, 18, 20]. Recently, GKPD proposed to re-parameterize the model weights using Kronecker-product decomposition and demonstrated promising performance [7]. PELA applies low-rank training to the pre-training phase to reduce the computational cost of training [6]. FLANC uses a customized tensor decomposition method to enable direct model aggregations across clients in the context of federated learning [23]. FedPara combines decomposition methods and the Hadamard product to enhance the representation capacity of re-parameterized models [12]. TKD-CPD jointly utilize Tucker and CP decompositions [25]. While these methods effectively reduce the number of trainable parameters and thus the computational and communication costs, they do not consider the inherent rank decline issue.
Rank Recovery Methods – Some regularization methods have been proposed to address the issue of inherent rank decline. Soft orthogonality (SO) is the basic regularization method which minimizes the difference between Gram matrix of the weight matrix and the identity matrix [32]. Double soft orthogonality (DSO) is a variant of SO which considers the regularization with overcomplete and undercomplete Gram matrices [1]. Spectral Restricted Isometry Property (SRIP) is a variant of RIP [2] that minimizes the spectral norm of the difference between the Gram matrix of a weight matrix and the identity matrix. Another study propose a sine-activated low-rank training strategy that is also designed to restore the model weight’s rank [15]. These methods commonly recover the effective rank of model weights, however, the rank cannot exceed the hard limit imposed by the low-rank reparameterization. In this study, we focus on how to overcome this limitation by adjusting the model rank at run-time.
@@ 3.1 Motivation
Recently, it has been theoretically shown that the rank of gradients tends to decrease as training progresses [11, 32]. Consequently, given a fixed training dataset and a large number of repeated training steps, the model weights are also expected to exhibit a similar reduction in rank.
@@
Therefore, running full-rank training during the middle of training is expected to mitigate the rapid decline in the effective rank of model weights. This insight motivates the design of a general model rank adjustment framework, which we describe in the following section.
@@ 4 Dynamic-rank Training Framework
In this section, we propose a general dynamic-rank training framework that temporarily increases the rank of the model’s weight matrices at selected points during training. We first formalize rank inflation and deflation, then discuss how to schedule rank adjustments. In particular, we provide two key insights for placing full-rank epochs within low-rank training to maximize rank restoration. Figure 2 illustrates the proposed framework.
@@
4.1 Model Rank Adjustment
We define two directions of model rank adjustment, inflation and deflation, as follows. For simplicity, we use the notation of SVD-style low-rank reparameterization, but it can be easily replaced with other methods such as Tucker or CP decompositions.
Model Inflation – Low-rank training can be implemented in two different forms, vanilla low-rank reparameterization and low-rank adaptation, LoRA [10]. Without loss in generality, we define model inflation process as follows. Given low-rank model weights $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ and base parameter $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$, the model is inflated such that $\mathbf{W}\leftarrow\mathbf{W}_{0}+\mathbf{A}\mathbf{B}^{\top}$. Note that we assume $k<r\leq n$. If the low-rank model follows the standard low-rank reparameterization, the initial weight matrix $\mathbf{W}_{0}$ is set to the zero matrix, i.e., $\mathbf{0}_{m\times n}$. Consequently, the maximum available rank is increased from $k$ to $n$. Please see Appendix for the case of convolutional layers.
Model Deflation – Given a model weight $\mathbf{W}$, the model is deflated by attaching a low-rank adaptor path next to the original weight such that $\mathbf{W}_{f}+\mathbf{A}\mathbf{B}^{\top}\leftarrow\mathbf{W}$, where $\mathbf{W}_{f}\in\mathbb{R}^{m\times n}$ is the given model weight matrix and $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are low-rank model weights. The provided weight matrix is frozen as $\mathbf{W}_{f}$ and $\mathbf{A}$ and $\mathbf{B}$ are trained instead. One can initialize $\mathbf{A}$ and $\mathbf{B}$ using either random distributions or zero matrices. Note that this naturally resembles LoRA when SVD is used as the reparameterization method.
These model rank adjustment procedures incur additional computations during training, however, when the epoch budget $E$ is sufficiently large, the extra computational cost typically becomes negligible. We now turn our attention to when the model rank should be inflated or deflated during training.
@@ 4.2 Model Rank Scheduling
Let the total training budget be $E$ epochs. We begin training in a low-rank form and increase the rank of model weights at epoch $I$. Then, at epoch $D$, we reduce the rank again. That is, training is performed in the full-rank space for DID-I epochs, and in the low-rank space for the remaining E(DI)E-(D-I) epochs. Under this setting, the key question is how to find good $I$ and $D$. Here, we present two empirical insights that motivate practical principles for selecting $I$ and $D$ to maximize rank recovery.
@@
Figure 3: CIFAR-10 (ResNet20) benchmark with various dynamic-rank training schedules. Inflate and Deflate indicate the epoch where the model rank is increased and decreased, respectively. The best accuracy is achieved when the full-rank epochs are located in both high-noise and low-noise regimes. By contrast, restricting full-rank epochs to only one of these regimes leaves the training loss insufficiently reduced.
@@
Full-rank training should be performed in both high-noise and low-noise regimes – We propose performing full-rank training both before and after the learning rate decay. This ensures that all ranks are sufficiently exposed to different learning rates during training. After deflating the model, only up to $k$ ranks are updated in the adaptor and the remaining nkn-k ranks remain unchanged in the frozen unless the model is inflated again. Thus, if the model is deflated before the learning rate decay and not inflated again in the rest of the training, the nkn-k ranks will not be trained under the low-noise regime. Those nkn-k ranks may lead to underfitting, resulting in lower model accuracy.
@@
In contrast, if full-rank training is performed only in the low-noise regime, the model is likely to suffer from poor generalization performance. It has been shown that training in a high-noise regime during the early phase improves generalization performance [22, 8, 21]. Therefore, if the model is inflated for the first time in the low-noise regime, the nkn-k ranks will be trained only with a small learning rate, potentially degrading the model’s generalization performance. Moreover, scheduling the full-rank epochs later reduces the training iterations for the adapter’s $k$ ranks, potentially causing underfitting.
@@
Figure 3 shows simple empirical results which validates our argument. We evaluated CIFAR-10 with ResNet-20 under three schedules that insert the full-rank training epochs at different points. The model is reparameterized with SVD, reducing its rank by half. Because the learning rate decays after epoch 100, the first and third settings correspond to full-rank epochs only in the high-noise and low-noise regimes, respectively. As expected, the best accuracy is achieved when full-rank epochs are located in both noise regimes. Note that the curves spike during deflation, as a randomly initialized low-rank adaptor is attached and begins training. The yellow and orange curves clearly show both underfitting and low accuracy. Based on this analysis, we conclude that full-rank training should be applied in both the high-noise and low-noise regimes to sufficiently train the model.
Full-rank training should be at the end of the high-noise regime and the early low-noise regime – We begin by analyzing the rank boundaries of the reconstructed model after deflation, as below.
@@
Table 2: Layer-wise comparison of singular-value spectral ratio ($\lambda$) across different rank-adjustment schedules. The $I$ and $D$ indicate rank inflation and deflation, respectively. ResNet20 was trained on CIFAR-10 for 150 epochs, with the learning rate reduced by a factor of 10 at epochs 100 and 130. Having full-rank epochs at the end of the high-noise regime yields the best results.
Proposition 1. Suppose model weight matrices: $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$, $\mathbf{A}\in\mathbb{R}^{m\times k}$, and $\mathbf{B}\in\mathbb{R}^{n\times k}$ and a re-constructed matrix $\mathbf{W}=\mathbf{W}_{0}+\mathbf{A}\mathbf{B}^{\top}$. Assuming that the effective rank of $\mathbf{W}_{0}$ is $r$, the rank of the reconstructed matrix $\mathbf{W}$ is bounded between rkr-k and $r+k$.

### v3

@@ Abstract
Low-rank training is a primary strategy for efficient deep learning, but it presents a fundamental challenge. It reduces computational cost, yet it permanently caps a model’s representational capacity and accelerates the rank collapse that diminishes its expressive power during training. We address this with dynamic-rank training, a framework built on the intuition that a model can temporarily escape its low-rank constraints to restore its full learning potential. Our approach strategically interleaves full-rank epochs within a low-rank schedule, with the timing of these restorative phases aligned with the learning rate’s noise regimes to maximize their effect. This enables the model to regain expressive power at critical stages of training by restoring the effective rank of its weights. Our extensive evaluations across various computer vision and natural language processing benchmarks show that the dynamic-rank method achieves the accuracy of full-rank models while retaining the computational advantages of low-rank training.
@@ 1 Introduction
Low-rank reparameterization methods have been actively studied for efficient training of large neural networks. Low-rank training strategies typically reduce the number of trainable parameters by applying matrix decompositions to a model’s weight matrices. While this approach lowers training cost, it permanently caps the maximum attainable rank of those matrices, thereby limiting the model’s ability to learn complex patterns. Moreover, recent studies report that the effective rank of the weights tends to decline during training (xie2017all, ; huang2025gradient, ). Therefore, to train large neural networks efficiently while preserving their learning capacity, it is crucial to address the decline in weight-matrix rank that occurs during low-rank training.
@@
Singular value decomposition(SVD)-based low-rank training (jaderberg2014speeding, ) is among the most popular re-parameterization techniques. Other tensor factorizations, such as Tucker (kim2015compression, ) and CP (lebedev2014speeding, ) decompositions, have likewise been adopted to enable low-rank training. More recently, low-rank fine-tuning methods have been proposed, including LoRA (hu2022lora, ), AdaLoRA (zhang2023adalora, ), LoRA-GA (wang2024loragalowrankadaptationgradient, ), DoRA (liu2024doraweightdecomposedlowrankadaptation, ), and SLTrain (han2024sltrainsparsepluslowrank, ). Although these approaches reduce the number of trainable parameters, they often overlook the progressive decline in the effective rank of the weights, which in turn compromises learning capability. Some regularization methods, such as soft orthogonal regularizer (xie2017all, ) and its variants (bansal2018can, ; kim2022revisiting, ), focus on tackling the rank decline issue. However, they incur higher computational cost and memory overhead, making them less practical for training large neural networks.
@@
This study explores how to mitigate the decline in the effective rank of model weights in low-rank training. Our key finding is that interleaving a few full-rank epochs within low-rank training effectively restores the model’s effective rank. Specifically, we analyze how run-time rank adjustment affects the singular value spectrum of the model weights, and we present a practical strategy for adjusting the rank during training to mitigate the decline in effective rank. Our theoretical analysis and empirical study demonstrate that dynamic rank adjustment matches the accuracy of full-rank training while retaining the system efficiency of low-rank methods.
@@
Another key finding is that the effectiveness of rank adjustment is closely coupled with the learning rate. Our study shows that full-rank epochs should be scheduled according to the noise scale induced by the learning rate. In general, the learning rate tends to be initialized to a large value and then progressively decayed during training. In this study, we analyze how the learning rate affects the gap between low-rank and full-rank updates and, based on this analysis, propose a general rank scheduling framework that maximizes the benefits of interleaving full-rank epochs within low-rank training. Table 1 provides a feature-wise comparison of various low-rank training methods.
@@
To the best of our knowledge, this is the first study to explore the benefits of interleaving full-rank training with low-rank training and to demonstrate the efficacy of a dynamic rank adjustment technique. We evaluate the proposed dynamic-rank training framework through extensive empirical studies on computer vision and natural language processing benchmarks. Furthermore, we benchmark our approach against state-of-the-art (SOTA) low-rank training techniques and regularization methods designed to restore effective rank. Across all experiments, we find that dynamic-rank training achieves accuracies comparable to full-rank training while significantly reducing computational costs, similar to conventional low-rank approaches. In addition, our results show that the proposed method integrates seamlessly with low-rank training when combined with soft-orthogonality (SO) regularization, confirming that the two techniques are complementary.
@@ 2 Background
Table 1: Feature-wise comparison across representative low-rank training methods.
Low-rank Training Method Parameter Rank Pre-training Noise-scale Decomposition
-efficient Recovery Compatible Aware Agnostic
SVD-based Low-rank (jaderberg2014speeding, )
TKD-CPD (phan2020stable, )
GKPD (hameed2022convolutional, )
Soft Orthogonality (xie2017all, )
SRIP+ (kim2022revisiting, )
LoRA (hu2022lora, )
AdaLoRA (zhang2023adalora, )
PELA (guo2024learning, )
SLTrain (han2024sltrainsparsepluslowrank, )
SparseLoRA (khaki2025sparseloraacceleratingllmfinetuning, )
Dynamic-rank (proposed)
Low-rank Re-parameterization – Low-rank reparameterization is a model approximation technique that is popularly used in large neural network training. Given a model weight matrix $\mathbf{W}$, a matrix decomposition method is applied to $\mathbf{W}$ before training begins. For example, if SVD is used, $\mathbf{W}\in\mathbb{R}^{m\times n}$ is decomposed to two smaller matrices $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ such that $\mathbf{W}=\mathbf{A}\mathbf{B}^{\top}$, where $k<m$ and $k<n$. As $k$ decreases, the total number of model parameters is reduced, thereby lowering the computational cost of training. However, the reconstructed weight matrix can have up to $k$ ranks, resulting in the limited learning capability.
Learning Rate and Noise Scale – In previous works, learning rate is known to play a key role in determining generalization performance of machine learning models. It has been theoretically shown that the noise scale $g$ in gradient approximation is determined by the learning rate and batch size, as follows (smith2017don, ): $g\approx\eta N/B$, where $\eta$ is learning rate, $N$ is the dataset size, and $B$ is the mini-batch size. This relation links learning rate schedules to the notion of noise scale. In early training, a large noise scale helps the model explore and shape the decision boundary (li2019towards, ; lee2023achieving, ), consistent with the common practice of decaying the learning rate. In this study, we investigate the relationship between the noise scale and the rank of model’s weights.
@@ 3 Related Work
Low-rank Reparameterization Methods – Several studies have explored the use of low-rank re-parameterization techniques including singular value decomposition, Tucker decomposition, and CP decomposition methods (jaderberg2014speeding, ; kim2015compression, ; lebedev2014speeding, ). Recently, GKPD proposed to re-parameterize the model weights using Kronecker-product decomposition and demonstrated promising performance (hameed2022convolutional, ). PELA applies low-rank training to the pre-training phase to reduce the computational cost of training (guo2024learning, ). FLANC uses a customized tensor decomposition method to enable direct model aggregations across clients in the context of federated learning (mei2022resource, ). FedPara combines decomposition methods and the Hadamard product to enhance the representation capacity of re-parameterized models (hyeon2021fedpara, ). TKD-CPD jointly utilize Tucker and CP decompositions (phan2020stable, ). While these methods effectively reduce the number of trainable parameters and thus the computational and communication costs, they do not consider the inherent rank decline issue.
Low-rank Fine-tuning Methods – Low-rank Adaptation (LoRA) (hu2022lora, ) is one of the most popular applications of re-parameterization technique, particularly in the context of fine-tuning. The importance of singular values under LoRA’s context is deeply analyzed in  (ke2024unveiling, ). AdaLoRA (zhang2023adalora, ) is a variant of LoRA which dynamically controls the rank of layers within a fixed total parameter budget. DoRA (liu2024doraweightdecomposedlowrankadaptation, ) separately fine-tunes the magnitude and the direction, based on LoRA. SLTrain (han2024sltrainsparsepluslowrank, ) leaves a sparse and frozen matrix besides the adaptor. LoRA-GA (wang2024loragalowrankadaptationgradient, ) initializes adapters using a subset of eigenvectors of gradient matrices. SparseLoRA (khaki2025sparseloraacceleratingllmfinetuning, ) decomposes pretrained weights using SVD.
Rank Recovery Methods – Some regularization methods have been proposed to address the issue of inherent rank decline. Soft orthogonality (SO) is the basic regularization method which minimizes the difference between Gram matrix of the weight matrix and the identity matrix (xie2017all, ). Double soft orthogonality (DSO) is a variant of SO which considers the regularization with overcomplete and undercomplete Gram matrices (bansal2018can, ). Spectral Restricted Isometry Property (SRIP) is a variant of RIP (candes2005decoding, ) that minimizes the spectral norm of the difference between the Gram matrix of a weight matrix and the identity matrix. Another study propose a sine-activated low-rank training strategy that is also designed to restore the model weight’s rank (ji2024efficient, ). These methods commonly recover the effective rank of model weights, however, the rank cannot exceed the hard limit imposed by the low-rank reparameterization. In this study, we focus on how to overcome this limitation by adjusting the model rank at run-time.
Figure 1: Comparison of the layer-wise singular-value spectral ratio ($\lambda$) across different model ranks during ResNet20 training on CIFAR-10. The left plot shows layer-wise $\lambda$ curves for full-rank training, while the right plot shows those for SVD-based low-rank training. We omit the legend since there are too many layers. Throughout the whole training, most layers in the re-parameterized model exhibit large $\lambda$ values, indicating convergence to a low-rank space.
@@ 3.1 Motivation
Recently, it has been theoretically shown that the rank of gradients tends to decrease as training progresses (huang2025gradient, ; xie2017all, ). Consequently, given a fixed training dataset and a large number of repeated training steps, the model weights are also expected to exhibit a similar reduction in rank.
@@
Therefore, having full-rank epochs in the middle of training is expected to mitigate the rapid decline in the effective rank of model weights. This insight motivates the design of a general model rank adjustment framework, which we describe in the following section.
@@ 4 Dynamic-rank Training Framework
In this section, we propose a general dynamic‑rank training framework that temporarily increases the rank of the model’s weight matrices at selected points during training. We first formalize rank inflation and deflation, then discuss how to schedule rank adjustments. In particular, we provide two key insights for placing full‑rank epochs within low‑rank training to maximize rank restoration. Figure 2 illustrates the proposed framework.
4.1 Model Rank Adjustment
We first define two directions of model rank adjustment, inflation and deflation, as follows. For simplicity, we use the notation of SVD-style low-rank reparameterization, but it can be easily replaced with other methods such as Tucker or CP decompositions.
@@ 4.1 Model Rank Adjustment
Model Inflation – Low-rank training can be implemented in two different forms, vanilla low-rank reparameterization and low-rank adaptation, LoRA (hu2022lora, ). Without loss in generality, we define model inflation process as follows. Given low-rank model weights $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ and base parameter $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$, the model is inflated such that $\mathbf{W}\leftarrow\mathbf{W}_{0}+\mathbf{A}\mathbf{B}^{\top}$. Consequently, the maximum available rank is increased from $k$ to $n$. If the low-rank model follows the standard low-rank reparameterization, the initial weight matrix $\mathbf{W}_{0}$ is set to the zero matrix, i.e., $\mathbf{0}_{m\times n}$. Please see Appendix for the case of convolutional layers.
Model Deflation – Given a model weight $\mathbf{W}$, the model is deflated by attaching a low-rank adaptor path next to the original weight such that $\mathbf{W}_{f}+\mathbf{A}\mathbf{B}^{\top}\leftarrow\mathbf{W}$, where $\mathbf{W}_{f}\in\mathbb{R}^{m\times n}$ is the given model weight matrix and $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are low-rank model weights. The provided weight matrix is frozen as $\mathbf{W}_{f}$, and only $\mathbf{A}$ and $\mathbf{B}$ are trained. $\mathbf{A}$ and $\mathbf{B}$ can be initialized using either random distributions or zero matrices. Note that this naturally resembles LoRA when SVD is used as the reparameterization method.
These model rank adjustment procedures incur extra computations during training, however, when the epoch budget $E$ is sufficiently large, the extra computational cost typically becomes negligible. We now turn our attention to when the model rank should be inflated or deflated during training.
@@ 4.2 Model Rank Scheduling
Here, we discuss how to determine appropriate timings for adjusting the model rank, taking into account the common practice of using learning rate decay for noise control in modern deep learning. Let the total training budget be $E$ epochs. Training begins in a low-rank form, with the rank of the model weights increased at epoch $I$ and reduced back at epoch $D$. That is, training is performed in the full-rank space for DID-I epochs and in the low-rank space for the remaining E(DI)E-(D-I) epochs. Under this setting, the key question is what values of $I$ and $D$ are optimal. We first present two practical principles for choosing $I$ and $D$ to maximize rank recovery.
@@
Figure 3: CIFAR-10 (ResNet20) benchmark with various dynamic-rank schedules. Inflate and Deflate indicate the epoch where the model rank is increased and decreased, respectively. The best accuracy is achieved when the full-rank epochs are located in both high-noise and low-noise regimes.
@@
Full-rank training should be performed in both high-noise and low-noise regimes – We propose having full-rank epochs both before and after the learning rate decay. This ensures that all ranks are sufficiently exposed to different learning rates during training. After deflating the model, only $k$ ranks are updated in the adaptor and the remaining nkn-k ranks stay frozen until the model is inflated again. Thus, if the model is deflated before learning rate decay, the nkn-k ranks do not get trained under the low-noise regime, potentially causing underfitting and reduced accuracy.
@@
In contrast, if full-rank training is performed only in the low-noise regime, the model is likely to suffer from poor generalization performance. It has been shown that training in a high-noise regime during the early phase improves generalization performance (li2019towards, ; he2019control, ; lee2023achieving, ). Therefore, if the model is inflated for the first time in the low-noise regime, the nkn-k ranks will be trained only with a small learning rate, potentially degrading the model’s generalization performance. Moreover, scheduling the full-rank epochs later reduces the training iterations for the adapter’s $k$ ranks, potentially causing underfitting.
@@
Figure 3 presents empirical results supporting our argument. We trained ResNet20 on CIFAR-10 under three schedules, each with full-rank training epochs inserted at different points. The model is reparameterized with SVD to reduce its rank by half. Because the learning rate decays after epoch 100, the first and third schedules correspond to full-rank epochs in the high-noise and low-noise regimes, respectively. As expected, the best accuracy is achieved when full-rank epochs are located in both noise regimes. The curves spike during deflation as a randomly initialized low-rank adaptor is attached and begins training. The yellow and orange curves clearly exhibit underfitting and low accuracy. Based on this analysis, we conclude that full-rank training should be applied in both high-noise and low-noise regimes to sufficiently train the model.
Full-rank training should be at the end of the high-noise regime and the early low-noise regime – First, we define the effective rank of $\mathbf{W}$, denoted by $r$, as the number of singular values that are significantly greater than zero. We analyze rank bounds of the reconstructed model as follows.
Proposition 1. Suppose model weight matrices: $\mathbf{W}_{0}\in\mathbb{R}^{m\times n}$, $\mathbf{A}\in\mathbb{R}^{m\times k}$, and $\mathbf{B}\in\mathbb{R}^{n\times k}$ and a re-constructed matrix $\mathbf{W}=\mathbf{W}_{0}+\mathbf{A}\mathbf{B}^{\top}$. Assuming that the effective rank of $\mathbf{W}_{0}$ is $r$, the rank of the reconstructed matrix $\mathbf{W}$ is bounded between $\max\{0,r-k\}$ and $\min\{m,n,r+k\}$.
Proof.
First, re-write each matrix as a summation of rank-1 matrices such that $\mathbf{W}_{0}=\sum_{j=1}^{r}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}$ and $\mathbf{A}\mathbf{B}^{\top}=\sum_{j=1}^{k}\mathbf{a}_{j}\mathbf{b}_{j}^{\top}$. Consequently, $\mathbf{W}=\sum_{j=1}^{r}\mathbf{u}_{j}\mathbf{v}_{j}^{\top}+\sum_{j=1}^{k}\mathbf{a}_{j}\mathbf{b}_{j}^{\top}$. The minimum rank of $\mathbf{W}$ is rkr-k if the $k$ rank-1 matrices in $\mathbf{A}\mathbf{B}^{\top}$ eliminate $k$ rank-1 components of $\mathbf{W}_{0}$. On the other hand, the maximum rank of $\mathbf{W}$ is $r+k$ if the $k$ rank-1 matrices in $\mathbf{A}\mathbf{B}^{\top}$ are orthogonal to all $k$ components of $\mathbf{W}_{0}$. Therefore, considering appropriate caps, the rank of the reconstructed matrix $\mathbf{W}$ is bounded between $\max\{0,r-k\}$ and $\min\{m,n,r+k\}$. ∎
Table 2: Comparison of singular-value spectral ratio ($\lambda$) across different rank-adjustment schedules. The $I$ and $D$ indicate rank inflation and deflation, respectively. ResNet20 was trained on CIFAR-10 for 150 epochs, with the learning rate reduced by a factor of 10 at epochs 100 and 130.
@@
To connect theory and practice, we compare weight ranks under different rank schedules in Table 2. We perform CIFAR-10 (ResNet‑20) training with the learning rate decaying at epoch 100. The rightmost three columns show $\lambda$ for layers 0, 10, and 18 (near the input, middle, and output). In the conventional low-rank training (None), $\lambda_{0}$ and $\lambda_{18}$ are very large, indicating widely spread spectra and effective rank loss. As full‑rank training is applied later, the singular value spectra become narrow (smaller $\lambda$ values). These results demonstrate that the full‑rank budget should be placed toward the end of the high‑noise regime.
Now, let us discuss the impact of full-rank training in the low-noise regime. Once the learning rate decays, the rank of the model weights does not dramatically change due to the reduced magnitude of updates. Instead, what matters more is the extent to which low-rank training pushes the model away from the full-rank convergence point. The following proposition formalizes this relationship.
Proposition 2. As the learning rate decays, the gap between the full-rank model update and the low-rank model update is expected to decrease.